OpenStudy (anonymous):

Find the square root 0.01

OpenStudy (anonymous):

0.01=1*10^-2 sqrt(1*10^-2)=plus-minus1*10^-1 or 0.1

OpenStudy (anonymous):

\[\sqrt{.01}=\sqrt{\frac{1}{100}}=\frac{\sqrt{1}}{\sqrt{100}}=\frac{1}{10}\]